OpenStudy (anonymous):

Simplify 2x^2-5x+3 x^2-4 --------- divided by ------------ 4x^2-9 4x^2+14x+12

OpenStudy (adilalvi):

u mean 1st divided by 2nd

OpenStudy (anonymous):

yes

OpenStudy (adilalvi):

it is very lenthy question.

OpenStudy (anonymous):

@Kainui

OpenStudy (gorv):

u know factorization??

OpenStudy (gorv):

2x^2-5x+3=2x^2-3x-2x+3 =x(2x-3)-1(2x-3)=(2x-3)(x-1)

OpenStudy (anonymous):

no

OpenStudy (gorv):

4x^2-9=(2x)^2-3^2=(2x-3)(2x+3)

OpenStudy (gorv):

\[\frac{ (2x-3)(x-1) }{ (2x-3)(2x+3)}\]

OpenStudy (gorv):

\[\frac{ x-1 }{ 2x+3 }\]

OpenStudy (anonymous):

this is the simplified form?

OpenStudy (gorv):

no this is first part ....the part before divide

OpenStudy (gorv):

now term after divide

OpenStudy (gorv):

x^2-4=x^2-2^2=(x-2)(x+2)

OpenStudy (gorv):

4x^2+14x+12=4x^2+8x+6x+14

OpenStudy (gorv):

4x^2+8x+6x+12 =4x(x+2)+6(x+2)=(x+2)(4x+6)=(x+2)*2*(2x+3)

OpenStudy (gorv):

(x-2)(x+2)/(2*(x+2)(2x+3)=(x-2)/(2*(2x+3)

OpenStudy (gorv):

\[\frac{ x-1 }{ 2x+3 }\div \frac{ x-2 }{2( 2x+3 )}\]

OpenStudy (gorv):

\[\frac{ x-1 }{ 2x+3 }*\frac{ 2(2x+3 )}{ x-2}\]

OpenStudy (gorv):

\[\frac{ 2*(x-1) }{ x-2 }\]

OpenStudy (gorv):

simplified

OpenStudy (anonymous):

woah that was alot of steps but i understand now thank you so much for helping me :)

OpenStudy (gorv):

welcome :)